# GraphArena: Benchmarking Large Language Models on Graph Computational Problems

### v1

@@ 2.2 Task Design
Problem Generation: For each of the 10 tasks, we randomly sample 500 small and 500 large graphs to create two distinct subsets, yielding a total of 10,000 graphs. The size ranges for these graphs are tailored to the complexity of the tasks. For the Neighbor and Distance tasks, small graphs contain 4 to 19 nodes, while large graphs contain 20 to 50 nodes. For Component, Diameter, MCP, MIS, and MVC, small graphs have 4 to 14 nodes, and large graphs range from 15 to 30 nodes. For MCS, GED and TSP, small graphs have 4 to 9 nodes and large graphs range from 10 to 20 nodes.
After sampling graphs with the given size, we encode graph problem into text based on templates, a common approach in previous benchmarks [75, 17]. An example of this can be seen in the bottom left of Figure 1. The process begins with an introduction and definition of the task, followed by a randomly selected example with the correct answer to demonstrate the problem-solving process. Subsequently, the actual problem to be solved is listed, including detailed graph information and task requirements. While the graphs used may not appear large, their conversion into text often results in problems containing up to 6,000 tokens, posing a significant long-range challenge. Detailed statistics on graph sizes and problem lengths, along with examples for each task, are provided in the appendix.
2.3 Evaluation Framework
Unlike previous benchmarks that only verify responses as yes, no, or a numeric value, we develop a more detailed evaluation framework to prevent pattern-based guesses. For each task, LLMs are required to output the relevant graph component or the path that contributes to the final answer, as indicated by the green sections in each graph in Figure 1. Responses are categorized into four types through a checker: (1) Correct: The solution satisfies both specified task requirements; (2) Suboptimal: The solution satisfies the first requirement but not the second; (3) Hallucinatory: The solution is in the correct format but does not satisfy either requirement; (4) Missing: The LLM fails to provide a solution in the correct format.
For example, in the bottom middle of Figure 1, Deepseek-V2 correctly identifies the maximum clique of four authors. In contrast, GPT-4o finds a feasible but suboptimal solution with three nodes, while Llama3-8b produces a hallucinatory response, erroneously linking Claudiu and Fabien—a connection that does not exist. If we merely report numbers as in existing graph problem benchmarks, this error would be misclassified as a correct answer.
Metrics. Given the rarity of missing results in our experiments (less than 1% in most cases), our primary metrics focus on Accuracy (the proportion of correct answers), Feasibility (the proportion of both correct and sub-optimal answers), and Hallucination (the proportion of hallucinatory responses). We also introduce ranking-based metrics for additional comparison, particularly useful when most models yield suboptimal solutions. These metrics include Mean Reciprocal Rank (MRR), Top-1 Probability, and Top-3 Probability.
3 Experiments
Experimental Setup. In our experiments, we evaluated ten popular LLMs in GraphArena, encompassing both closed-source and open-source models. This included single models and mix-of-experts with diverse scales. Among the closed-source models, we tested the latest GPT-4o (gpt-4o-2024-05-13) [54], the well-known GPT-3.5 (gpt-3.5-turbo-0125) [56], and the cost-effective Claude3-haiku [6]. In the open-source category, our assessments included the Llama series (Llama3-70b-Instruct and Llama3-8b-Instruct) from Meta, the Qwen series (Qwen1.5-72b-Chat and Qwen1.5-8b-Chat) from AliCloud, and Gemma-7b from Google Cloud [1]. Additionally, we evaluated two mix-of-expert LLMs: Deepseek-V2 [10] with 230 billion parameters (21 billion active) and Mixtral-7x8b [35] featuring 47 billion parameters (13 billion active).
For closed-source LLMs and open-source LLMs larger than 10 billion parameters, we utilized services from cloud providers, including OpenAI, Claude, AliCloud [23], Deepseek [60], and AIMLAPI [7]. For smaller models with fewer than 10 billion parameters, inferences were conducted on our local machines equipped with four NVIDIA H800 PCIe 80GB GPUs. EGiven the high costs and time-intensive nature of evaluating LLMs, we limited our runs to a single instance per model across 10,000 problems in GraphArena. To encourage consistency in responses, we used a low temperature setting of 0.1 and imposed no constraints on the output length of any model. All responses were documented for public reference.
3.1 Main Results
Polynomial-Time Tasks (Small) Polynomial-Time Tasks (Large)
LLM MRR Top-1 Top-3 Acc. Fea. Hallu. MRR Top-1 Top-3 Acc. Fea. Hallu.
GPT-4o 0.82 78.00 81.53 71.90 82.30 17.45 0.63 55.95 63.29 40.65 59.00 39.55
GPT-3.5 0.46 36.58 42.66 34.00 52.75 47.25 0.30 18.37 25.82 14.20 30.30 69.55
Claude-haiku 0.53 45.51 49.08 43.10 53.90 46.00 0.34 23.89 29.63 18.85 33.40 66.50
Llama3-70b 0.72 66.42 70.80 61.20 72.70 27.30 0.54 45.58 53.93 31.55 50.95 49.05
Llama3-8b 0.40 30.74 34.98 28.55 45.30 53.90 0.23 12.45 15.85 9.40 19.65 78.20
Qwen1.5-72b 0.50 41.62 46.87 39.05 58.05 41.80 0.34 24.51 28.14 20.45 28.85 70.80
Qwen1.5-7b 0.23 12.79 14.67 11.85 20.75 79.10 0.16 6.11 7.27 4.60 8.40 91.20
Deepseek-V2 0.63 56.11 62.53 51.35 64.65 34.95 0.44 34.50 41.59 24.70 38.00 61.15
Mixtral-7x8b 0.48 39.81 45.04 37.80 53.00 46.35 0.31 19.24 28.00 13.85 31.75 67.20
Gemma-7b 0.35 26.29 29.39 25.15 38.60 61.40 0.20 10.31 12.21 9.20 16.05 83.95
NP-Complete Tasks (Small) NP-Complete Tasks (Large)
GPT-4o 0.62 52.36 62.84 40.83 74.20 23.23 0.42 28.68 43.28 5.93 49.17 45.60
GPT-3.5 0.47 33.28 46.91 25.17 73.20 26.33 0.26 11.32 22.81 2.27 38.57 59.77
Claude-haiku 0.56 43.50 58.09 33.47 78.00 21.10 0.33 17.98 34.67 3.93 40.90 58.13
Llama3-70b 0.58 47.50 60.47 36.77 75.47 24.53 0.37 21.01 40.22 4.80 48.77 50.07
Llama3-8b 0.37 25.15 33.40 20.17 58.07 41.73 0.22 7.11 21.39 1.37 31.17 68.77
Qwen1.5-72b 0.48 36.25 48.14 28.30 69.73 27.40 0.30 13.62 31.67 3.93 47.97 51.63
Qwen1.5-7b 0.28 16.00 22.67 12.70 50.17 49.63 0.18 4.85 12.43 1.13 25.63 73.83
Deepseek-V2 0.56 43.48 58.68 33.70 74.17 24.47 0.35 19.80 36.68 5.43 45.00 54.97
Mixtral-7x8b 0.33 19.76 28.15 15.57 59.43 37.43 0.21 7.99 15.52 2.13 24.20 75.50
Gemma-7b 0.31 16.95 26.69 12.93 57.13 42.60 0.21 7.17 16.20 1.67 29.40 64.63
In Table 1, we present the average performance of ten LLMs across four polynomial-time and six NP-complete tasks, assessed on both small and large graphs. As expected, GPT-4o consistently ranks as the top performer across most metrics and settings, with Llama3-70b closely following as the premier open-source model. Performance generally declines for all LLMs as graph sizes increase and tasks transition from polynomial-time to NP-complete, expanding the search space significantly. GPT-4o, despite its capabilities, struggles notably with NP-complete tasks on large graphs, achieving only a 5.93% accuracy rate. It indicates the most challenging scenario in GraphArena is still much larger than current LLMs’ upper bound.
GraphArena also reveals a significant performance gap between LLMs with different parameter numbers. For example, in polynomial-time tasks, the disparity between Llama3-70b and Llama3-8b is notable—61.2% vs. 28.6% on small graphs and 31.6% vs. 9.4% on large graphs. This gap is significantly wider than in other benchmarks, such as GSM8K (93.0% vs. 79.6%) and GPQA (39.5% vs. 34.2%). A similar trend is observed in the Qwen series. Moreover, models with fewer parameters are more prone to hallucination, with ratios of 78.2% for Llama3-8b, 91.2% for Qwen1.5-7b, and 84.0% for Gemma-7b, indicating that GraphArena’s challenges necessitate more advanced reasoning capabilities that larger LLMs are better equipped to manage.
To closely examine model performance on individual tasks, we analyzed the feasibility and accuracy of five selected LLMs—Claude3-haiku, Deepseek-V2, GPT-4o, Llama3-70b, and Mixtral-7x8b—as shown in Figure 2. Results for the remaining five LLMs are demonstrated in Figure 5 in the appendix. The general trends align with Table 1, showing decreased performance in larger graphs and more complex tasks. A detailed analysis shows that model performance varies substantially across different tasks. For instance, tasks such as Diameter, MCS, and MVC demonstrate notably low feasibility. This is attributed to their complex verification processes: verifying Diameter requires solving a shortest path problem, while verifying MCS involves graph isomorphism. Regarding accuracy, NP-complete tasks generally perform worse than polynomial-time tasks. Notably, the Diameter task shows the poorest performance among polynomial-time tasks due to its demand for a comprehensive global search. This indicates that task complexity significantly impacts accuracy.
3.2 In-depth Analysis of Hallucination
This section delves into the phenomenon of hallucination in LLMs, specifically aiming to address two critical questions: (1) What are the underlying causes of hallucination in GraphArena? (2) Are existing strategies able to mitigate the issue?
Our primary findings suggest that the hallucination ratio is linked with the number of model parameters, the complexity of its feasibility checks, and the scale of the graphs. To further illustrate this, we have plotted the hallucination probability against varying node sizes for three selected tasks in Figure 3. Additional data for seven other tasks are detailed in Figures 6 and 7 in the appendix. These plots reveal a significant, almost linear increase in hallucination probability as node size grows from 5 to 30. Notably, for GPT-4o on the Diameter task, the hallucination probability rises from 18% with a node size of 5 to 80% with a node size of 30, underscoring that larger graph sizes and their associated challenges in long text and exploding search space are major contributors to hallucination.
Chain-of-Thought (CoT) Prompting. Several strategies have been proposed to mitigate hallucinations in LLMs [9, 34, 87, 77]. We explore the effectiveness of a representative approach, CoT prompting, which encourages the model to generate intermediate reasoning steps towards a conclusion [78, 26, 52]. We design manually crafted examples for the Graph Diameter and Connected Component tasks, assessing the impact of using zero to four CoT examples. As shown in figure 4, a general enhancement in model accuracy is observed. Specifically, the performance of Deepseek-V2 notably improves after incorporating just one CoT example for the Component task. Nonetheless, Claude3-haiku and Mixtral-7x8b exhibit only marginal improvements. This suggests that, while CoT prompting contributes to reducing errors, it remains insufficient to fully resolve hallucinations.
Graph-aware LLMs. Several methods have been developed to fine-tune LLMs using graph-specific datasets, aiming to create graph-aware LLMs capable of understanding graph structures and solving basic graph tasks. To investigate whether this can help alleviate the issue of hallucinations, we test GraphArena on GraphWiz [17], an open-source model derived from Llama2-7b and fine-tuned using Direct Preference Optimization [61] across nine graph problems. While GraphWiz shows proficiency on the tasks it was initially trained on, it struggles on the GraphArena benchmark, with accuracies only ranging from 0.2% to 1.2% across four small-scale polynomial-time tasks. This can be attributed primarily to two factors: the absence of task overlap between GraphArena and the GraphWiz training corpus, and the model’s tendency to rely on pattern recognition rather than problem understanding.
@@ 3.2 In-depth Analysis of Hallucination
Table 2: Performance comparison of LLMs and classic solvers on TSP for both small and large graphs. The table shows the percentage of problems where LLMs wins, ties, or loses against three classic solvers: Random, Greedy, and Christofides.
Travelling Salesman Problem (Small Graphs)
Solver Random Greedy Christofides
LLM win tie lose win tie lose win tie lose
GPT-4o 67.4 16.2 16.4 20.6 43.4 36.0 15.4 33.0 51.6
GPT-3.5 58.4 13.6 28.0 13.0 30.2 56.8 7.4 24.0 68.6
Claude3-haiku 54.8 14.4 30.8 15.4 25.2 59.4 9.4 21.0 69.6
Llama3-70b 59.8 18.4 21.8 12.0 36.6 51.4 8.4 23.8 67.8
Llama3-8b 43.0 13.6 43.4 12.2 15.4 72.4 6.4 15.6 78.0
Qwen1.5-72b 49.4 16.0 34.6 13.2 25.0 61.8 7.0 22.8 70.2
Qwen1.5-7b 33.6 11.2 55.2 10.8 14.2 75.0 6.0 14.8 79.2
Deepseek-V2 71.0 15.4 13.6 21.0 42.0 37.0 15.8 32.2 52.0
Mixtral-7x8b 43.6 16.8 39.6 10.6 22.6 66.8 6.0 18.0 76.0
Gemma-7b 27.6 15.0 57.4 8.8 13.8 77.4 5.0 13.0 82.0
Travelling Salesman Problem (Large Graphs)
GPT-4o 82.6 0.0 17.4 9.2 2.8 88.0 5.4 0.4 94.2
GPT-3.5 66.8 0.0 33.2 0.4 0.0 99.6 0.2 0.0 99.8
Claude3-haiku 72.6 0.0 27.4 1.0 0.0 99.0 0.6 0.0 99.4
Llama3-70b 79.6 0.0 20.4 1.4 0.0 98.6 0.6 0.0 99.4
Llama3-8b 29.8 0.0 70.2 0.2 0.0 99.8 0.0 0.0 100.0
Qwen1.5-72b 26.4 0.0 73.6 0.0 0.0 100.0 0.0 0.0 100.0
Qwen1.5-7b 23.0 0.0 77.0 0.0 0.0 100.0 0.0 0.0 100.0
Deepseek-V2 67.2 0.0 32.8 4.4 0.2 95.4 2.2 0.0 97.8
Mixtral-7x8b 23.6 0.0 76.4 0.0 0.0 100.0 0.0 0.0 100.0
Gemma-7b 3.8 0.0 96.2 0.0 0.0 100.0 0.0 0.0 100.0
3.3 Comparison with Classic Solvers
To better understand the performance of LLMs on NP-complete tasks, we compare them with several classic solvers on the Travelling Salesman Problem (TSP) task. We consider the following three solvers: (1) Random: randomly generates a feasible solution, (2) Greedy: selects a node at each step that is not already in the cycle and whose connection to the previous node adds the least cost to the cycle, and (3) Christofides: computes a 3/2-approximation of the TSP in a complete undirected graph using the Christofides algorithm [21], which is considered a better heuristic algorithm compared to Greedy. For each solver, the percentage of problems where the LLM wins, ties, or loses against the classic solvers is reported in Table 2.
In the context of small graphs, GPT-4o demonstrates a strong performance with a win rate of 67.4% against the Random solver, 20.6% against the Greedy solver, and 15.4% against the Christofides solver. Other models such as GPT-3.5 and Claude3-haiku show competitive performance but with slightly lower win rates. Smaller Models like Llama3-8b and Gemma-7b exhibit higher loss rates, indicating less effectiveness in solving TSP compared to the classic solvers.
For large graphs, GPT-4o has an even better win rate of 82.6% against the Random solver, but its performance drops significantly against the Greedy and Christofides solvers. Most other LLMs show similar trends. However, GPT-4o can still win in 27 problems compared to the Christofides solver, highlighting the potential of LLMs to provide strong heuristics in combinatorial optimization tasks.
4 Related Work
LLM Evaluation. Early evaluations of text generation focused on fundamental skills, including similarity measures such as ROUGE [46], METEOR [8], and BLEU [58] for machine translation [47, 64]. They also included inherent scores like Diversity [41] and Perplexity [33] for open-domain conversations [86, 43], and factual measures such as F-score and exact match for machine reading comprehension [62, 14] and question answering [90, 38]. Recently, numerous benchmarks have employed more advanced tests to assess the high-level capabilities of LLMs [30, 22, 50, 48, 15, 40, 74, 16, 71]. For instance, the Massive Multitask Language Understanding benchmark [29] covers 57 subjects across STEM, the humanities, the social sciences, and more. The Beyond the Imitation Game Benchmark [65] includes over 200 tasks, ranging from language understanding to interaction and behavior. Human preference remains a critical aspect of LLM evaluation [88, 42, 20]. Chatbot Arena [20] employs a pairwise comparison approach and leverages input from a diverse user base on crowd-sourcing. Our proposed GraphArena benchmark aims to evaluate a wide range of reasoning capabilities of LLMs on polynomial-time and NP-complete graph Computational problems, offering a valuable supplement to the existing benchmarks.
LLM on Graphs. The synergy between graphs and LLMs has sparked considerable interest due to their bi-directional benefits: integrating graph-based approaches can enhance the reasoning abilities of LLMs, enabling them to tackle complex logical tasks such as mathematical problem-solving [85], code generation [28], and knowledge reasoning [83, 57], among others. Conversely, LLMs offer powerful capabilities to augment graph analytics and learning, particularly for predictive tasks in text-attributed networks [25, 75, 18, 19, 70]. Compared with machine learning tasks on graphs such as node classification and anomaly detection [68, 69], graph computational problems pose more challenges for LLMs as they require a deeper understanding of structural information and long-range multi-step reasoning [51, 76, 67, 12]. GraphQA [25] and NLGraph [75] are two early benchmarks in this domain, focusing on basic graph problems using small-scale synthetic graphs. VisionGraph [45] and GITA [79] introduce multimodal graph reasoning tasks, which require LLMs to reason over both text and image modalities related to the graph. GraphWiz [17] introduces an instruction-tuning dataset designed to equip language models with the ability to tackle graph problems using explicit reasoning paths.
5 Conclusion
This paper introduces GraphArena, a benchmarking tool designed to evaluate the reasoning capabilities of LLMs through graph computational challenges. GraphArena features a realistic graph collection, carefully selected tasks, and a rigorous evaluation framework. Our evaluation of ten LLMs across 10,000 GraphArena problems reveals persistent hallucination issues, especially with larger graphs, more complex tasks, and models with fewer parameters. Despite implementing advanced strategies such as chain-of-thought prompting and domain-specific fine-tuning, these issues remain unresolved. Future efforts include investigating LLMs’ ability to write code, integrating multi-modality to better address these challenges, and exploring new techniques to mitigate hallucinations and improve reasoning capabilities.
Appendix A Additional Experiments
We provide detailed statistics and examples to further explain the tasks and performance evaluations discussed in the main text. Table 3 presents comprehensive statistics for each task in GraphArena, including the average and maximum number of nodes edges, and text length counted by number of characters. Figure 5 offers a feasibility and accuracy comparison of the remaining five LLMs on each individual task, complementing the analysis in Figure 2. Figures 6 and 7 illustrate the influence of graph size on hallucination probability for the remaining seven tasks, as the complement of Figure 3. Tables LABEL:table:COT_Component_prompt through LABEL:table:TSP_prompt provide detailed examples of prompts for each task. For the Connected Component and Graph Diameter tasks, chain-of-thought (CoT) prompting with step-by-step solution demonstrations is provided.
Table 3: Statistics of problems in each task of GraphArena, including the average and maximum number of nodes ($V$), edges ($E$), and text length ($T$).
Problems avg. $\left|V\right|$ max. $\left|V\right|$ avg. $\left|E\right|$ max. $\left|E\right|$ avg. $\left|T\right|$ max. $\left|T\right|$
Common Neighbor 19.4 50 68.6 409 4323.5 18014
Shortest Distance 19.6 50 36.8 209 5639.5 23731
Connected Component 13.5 30 14.4 122 1785.4 5503
Graph Diameter 13.4 30 22.0 128 3741.3 11783
Maximum Clique Problem 13.2 30 39.1 154 3025.0 7921
Maximum Independent Set 12.8 30 19.5 141 1902.8 6182
Minimum Vertex Cover 14.1 30 23.3 151 2178.4 6429
Maximum Independent Set 9.2 20 8.9 22 1025.1 1187
Graph Edit Distance 9.4 20 9.2 23 1643.7 2142
Travelling Salesman Problem 9.5 20 50.6 190 1806.1 4741
@@ Appendix A Additional Experiments
Table 13: An example of the Travelling Salesman Problem.

### v2

@@ Appendix B Additional Task Information
Table 4 presents comprehensive statistics for each task in GraphArena, including the average and maximum number of nodes, edges, and text length (in characters). Tables LABEL:table:COT_Component_prompt-LABEL:table:TSP_prompt provide detailed examples of prompts for each task. For the Connected Component and Graph Diameter tasks, chain-of-thought prompting with step-by-step solution demonstrations is provided.
Appendix C Additional Experimental Results
Figure 8 offers a feasibility and accuracy comparison of the remaining five LLMs on individual tasks, complementing the analysis in Figure 2. Figures 9 and 10 illustrate the influence of graph size on hallucination probability for the remaining seven tasks, complementing Figure 4. Table 5 extends the analysis in Figure 3, showing the percentage of problems where each of the 10 LLMs wins, ties, or loses against Random, Greedy, and Approximated graph algorithms.
Comparison with graph neural networks (GNNs). Directly comparing GNNs and LLMs in GraphArena is challenging due to their fundamentally different paradigms. LLMs are multi-task, unsupervised models that process natural language inputs, while GNNs are typically task-specific, supervised models that operate on network data. For simplicity, we configured GNNs to directly predict the answer rather than generating the solution path or component. Although it is feasible to let GNNs generate detailed solutions, different tasks like GED and MCS would require specific architectural modifications. We compared three representative GNNs: Graph Isomorphism Network (GIN) Xu et al. (2019), Graph Attention Network (GAT) Veličković et al. (2017), and Graph Sample and Aggregate (GSAGE) Hamilton et al. (2017). We generated an equivalent amount of training data for GNNs using the same framework as GraphArena. Table 6 shows the average accuracy for these GNNs on Polynomial-time and NP-complete problems across small and large graphs. The results indicate that Claude-3.5-sonnet demonstrates superior performance in most scenarios. While GNNs achieve the best results on large-scale NP-complete problems, recent studies have theoretically proved the limitations in GNNs’ computational capabilities Zhang et al. (2024a). For instance, message passing neural networks cannot even count triangles. Given the limitations, we posit that GNNs might be performing pattern-based regression rather than genuinely solving graph problems. In contrast, LLMs appear more aligned with true problem-solving.
Comparison of different graph tokenizers. In Table 7, we examined the performance of three graph tokenizers—Edge List, Adjacency List, and Adjacency Matrix—on GPT-4o across two selected tasks: Shortest Distance and TSP, both on small and large graphs. As shown in The Edge List tokenizer demonstrated relatively stable performance across various tasks, with feasibility and accuracy scores consistently higher than those of the Adjacency Matrix, particularly in the Shortest Distance tasks where it achieved 0.914 and 0.902 accuracy for small graphs, and 0.788 and 0.720 for large graphs. Although the Adjacency List tokenizer slightly outperformed the Edge List in some instances, such as in the TSP (Large) task with a lower hallucination rate of 0.082, the Edge List remains the most stable option. Combined with its widespread use in prior studies like NLGraph Wang et al. (2023) and GraphWiz Chen et al. (2024a), we use Edge List as our primary graph tokenizer for the main experiments, as shown in Tables LABEL:table:COT_Component_prompt to LABEL:table:TSP_prompt. Conversely, the Adjacency Matrix often introduced redundant information, particularly in sparse graphs, leading to decreased performance, as evidenced by its 0.436 hallucination rate and 0.448 accuracy in the Shortest Distance (Large) task.
Comparison of Real-World and Synthetic Graphs. Table 8 presents the evaluation of GPT-4o on two polynomial-time tasks: Common Neighbor and Shortest Distance. Our findings indicate that GPT-4o achieves notably higher accuracy and lower hallucination rates on synthetic Erdős-Rényi graphs compared to real-world graphs of similar size. This disparity in performance suggests that synthetic graphs may fail to adequately represent the complex topological features and irregular structures found in real-world networks. Consequently, relying exclusively on synthetic benchmarks could lead to an overestimation of a model’s generalization capabilities. This observation motivates us to incorporate samples from real-world networks into GraphArena.
Table 4: Summary of task statistics in GraphArena, detailing the average and maximum values for nodes ($V$), edges ($E$), text length ($T$) in characters, graph density ($D$), and edge connectivity ($EC$) across problem instances.
Problems $\overline{\left|V\right|}$ $\left|V\right|_{\text{max}}$ $\overline{\left|E\right|}$ $\left|E\right|_{\text{max}}$ $\overline{\left|T\right|}$ $\left|T\right|_{\text{max}}$ $\overline{D}$ $\overline{EC}$
Common Neighbor 19.4 50 68.6 409 4323.5 18014 0.44 1.37
Shortest Distance 19.6 50 36.8 209 5639.5 23731 0.28 0.61
Connected Component 13.5 30 14.4 122 1785.4 5503 0.23 0.26
Graph Diameter 13.4 30 22.0 128 3741.3 11783 0.31 1.16
Maximum Clique Problem 13.2 30 39.1 154 3025.0 7921 0.52 1.70
Maximum Independent Set 12.8 30 19.5 141 1902.8 6182 0.30 1.08
Minimum Vertex Cover 14.1 30 23.3 151 2178.4 6429 0.31 1.17
Maximum Common Subgraph 9.2 20 8.9 22 1025.1 1187 0.30 1.04
Graph Edit Distance 9.4 20 9.2 23 1643.7 2142 0.29 1.03
Traveling Salesman Problem 9.5 20 50.6 190 1806.1 4741 1.00 8.46
@@ Appendix C Additional Experimental Results
Table 5: The percentage of problems where LLM wins, ties, or loses against three types of classic graph algorithms: Random, Greedy, and Approximated.
Travelling Salesman Problem (Small) Travelling Salesman Problem (Large)
Algorithm Random Greedy Approximated Random Greedy Approximated
LLM win/tie/lose win/tie/lose win/tie/lose win/tie/lose win/tie/lose win/tie/lose
GPT-4o 76.2/15.2/8.6 23.6/49.0/27.4 18.8/36.2/45.0 95.2/0.0/4.8 18.8/4.2/77.0 10.6/0.6/88.8
GPT-4o-mini 67.6/14.8/17.6 17.2/40.4/42.4 12.4/29.4/58.2 70.6/0.0/29.4 0.6/0.0/99.4 0.0/0.0/100.0
Claude-3.5-sonnet 77.4/15.8/6.8 24.6/52.2/23.2 17.4/40.6/42.0 99.6/0.0/0.4 33.2/8.4/58.4 20.8/1.0/78.2
Llama3-70b 65.0/13.4/21.6 12.0/36.6/51.4 8.4/23.8/67.8 78.6/0.0/21.4 1.4/0.0/98.6 0.6/0.0/99.4
Llama3-8b 47.0/13.6/39.4 12.2/15.4/72.4 6.4/15.6/78.0 30.6/0.0/69.4 0.2/0.0/99.8 0.0/0.0/100.0
Qwen2.5-72b 45.6/11.6/42.8 14.6/19.0/66.4 9.0/19.6/71.4 30.0/0.0/70.0 1.0/0.4/98.6 0.8/0.0/99.2
Deepseek-v2 72.6/15.4/12.0 21.0/42.0/37.0 15.8/32.2/52.0 66.8/0.0/33.2 4.4/0.2/95.4 2.2/0.0/97.8
Mixtral-7x8b 46.4/14.2/39.4 10.6/22.6/66.8 6.0/18.0/76.0 24.4/0.0/75.6 0.0/0.0/100.0 0.0/0.0/100.0
Gemma-7b 30.2/12.6/57.2 8.8/13.8/77.4 5.0/13.0/82.0 3.4/0.0/96.6 0.0/0.0/100.0 0.0/0.0/100.0
GLM-4-plus 69.8/14.0/16.2 19.6/39.4/41.0 13.2/31.0/55.8 84.0/0.0/16.0 5.2/0.2/94.6 2.6/0.4/97.0
Graph Edit Distance (Small) Graph Edit Distance (Large)
GPT-4o 71.0/14.6/14.4 35.0/49.2/15.8 12.4/30.8/56.8 94.6/1.4/4.0 49.6/18.2/32.2 34.2/3.8/62.0
GPT-4o-mini 57.2/12.2/30.6 30.2/37.8/32.0 10.6/28.4/61.0 47.0/0.8/52.2 29.4/10.2/60.4 25.4/3.0/71.6
Claude-3.5-sonnet 72.2/12.6/15.2 41.0/43.2/15.8 13.0/33.6/53.4 82.4/0.8/16.8 52.4/13.2/34.4 35.6/5.4/59.0
Llama3-70b 69.6/14.8/15.6 41.2/30.0/28.8 10.6/32.0/57.4 80.0/0.8/19.2 41.8/10.6/47.6 28.4/3.4/68.2
Llama3-8b 14.2/5.2/80.6 10.8/6.2/83.0 3.0/7.2/89.8 11.8/0.0/88.2 10.8/0.4/88.8 6.6/1.0/92.4
Qwen2.5-72b 64.0/10.0/26.0 37.4/33.2/29.4 11.2/34.6/54.2 20.8/0.2/79.0 14.0/3.2/82.8 10.8/1.8/87.4
Deepseek-v2 61.6/14.0/24.4 35.4/32.8/31.8 10.6/26.8/62.6 82.4/1.2/16.4 45.6/13.2/41.2 32.6/4.4/63.0
Mixtral-7x8b 37.8/12.4/49.8 17.6/21.4/61.0 6.0/11.4/82.6 43.6/1.2/55.2 15.8/4.4/79.8 9.4/1.2/89.4
Gemma-7b 53.6/22.4/24.0 30.8/21.0/48.2 3.6/6.4/90.0 75.4/1.2/23.4 43.0/7.2/49.8 25.0/3.4/71.6
GLM-4-plus 72.2/16.0/11.8 38.0/46.8/15.2 14.2/28.6/57.2 94.0/0.8/5.2 52.0/21.6/26.4 37.6/3.8/58.6
Maximum Independent Set (Small) Maximum Independent Set (Large)
GPT-4o 67.6/12.4/20.0 0.0/63.8/36.2 2.2/64.8/33.0 29.8/1.0/69.2 0.0/5.4/94.6 0.4/8.2/91.4
GPT-4o-mini 63.8/11.0/25.2 0.0/42.8/57.2 0.8/44.6/54.6 40.4/3.8/55.8 0.0/1.8/98.2 0.0/2.4/97.6
Claude-3.5-sonnet 61.8/11.4/26.8 0.0/68.0/32.0 1.8/68.0/30.2 25.2/1.0/73.8 0.0/7.8/92.2 0.8/8.8/90.4
Llama3-70b 62.4/21.4/16.2 0.0/36.8/63.2 0.6/38.2/61.2 42.6/3.4/54.0 0.2/2.6/97.2 0.6/3.8/95.6
Llama3-8b 46.0/10.0/44.0 0.4/28.0/71.6 1.4/29.2/69.4 14.2/0.4/85.4 0.0/1.0/99.0 0.2/1.4/98.4
Qwen2.5-72b 52.0/21.8/26.2 0.0/3.8/96.2 0.6/5.0/94.4 39.0/4.2/56.8 0.0/1.0/99.0 0.4/1.8/97.8
Deepseek-v2 55.4/15.6/29.0 0.0/36.2/63.8 0.8/37.4/61.8 12.2/0.6/87.2 0.0/3.2/96.8 0.4/3.6/96.0
Mixtral-7x8b 44.0/16.2/39.8 0.0/11.6/88.4 0.4/12.6/87.0 28.4/2.8/68.8 0.2/0.8/99.0 0.6/0.6/98.8
Gemma-7b 26.6/5.6/67.8 0.0/21.2/78.8 0.0/21.2/78.8 6.0/0.0/94.0 0.0/0.6/99.4 0.0/0.6/99.4
GLM-4-plus 64.2/11.6/24.2 0.2/57.2/42.6 2.2/58.8/39.0 34.4/0.4/65.2 0.0/3.8/96.2 0.6/5.4/94.0
Minimum Vertex Cover (Small) Minimum Vertex Cover (Large)
GPT-4o 40.6/7.6/51.8 39.8/7.0/53.2 30.0/16.4/53.6 19.0/1.2/79.8 13.6/2.6/83.8 12.0/3.8/84.2
GPT-4o-mini 34.6/3.2/62.2 27.4/8.2/64.4 25.4/9.6/65.0 21.6/0.2/78.2 11.8/3.6/84.6 11.4/6.2/82.4
Claude-3.5-sonnet 16.6/1.2/82.2 16.6/1.2/82.2 12.8/5.0/82.2 8.6/0.0/91.4 8.0/0.6/91.4 7.4/1.0/91.6
Llama3-70b 41.0/6.0/53.0 40.0/7.4/52.6 32.0/14.4/53.6 17.0/0.0/83.0 12.2/4.4/83.4 12.0/3.8/84.2
Llama3-8b 38.8/9.6/51.6 32.4/11.8/55.8 27.0/15.8/57.2 17.6/0.0/82.4 10.2/2.0/87.8 9.4/4.2/86.4
Qwen2.5-72b 2.8/0.8/96.4 3.4/0.2/96.4 2.0/1.6/96.4 1.6/0.0/98.4 0.6/0.0/99.4 0.6/0.0/99.4
Deepseek-v2 45.8/9.2/45.0 36.6/14.8/48.6 31.6/17.0/51.4 29.0/1.8/69.2 10.6/8.4/81.0 11.2/8.8/80.0
Mixtral-7x8b 28.2/11.4/60.4 15.4/20.4/64.2 12.6/19.0/68.4 21.2/0.8/78.0 6.2/3.8/90.0 5.6/4.4/90.0
Gemma-7b 38.2/14.2/47.6 15.0/30.4/54.6 15.0/25.6/59.4 15.8/0.2/84.0 4.8/2.4/92.8 3.6/4.4/92.0
GLM-4-plus 53.0/10.4/36.6 31.2/20.4/48.4 29.0/23.6/47.4 33.8/1.6/64.6 18.4/7.2/74.4 15.4/9.4/75.2
Table 6: Average accuracy (%) for three representative GNNs on Polynomial-time and NP-complete problems across small and large graphs.
Models P (Small) P (Large) NP (Small) NP (Large)
GIN 59.9 39.6 33.8 25.8
GAT 42.3 22.5 32.3 17.9
GSAGE 41.0 22.6 33.7 17.6
Claude-3.5-sonnet 82.2 58.7 47.8 7.2
Table 7: Performance comparison for different graph tokenizers on GPT-4o across tasks.
Tasks Distance (Small) Distance (Large) TSP (Small) TSP (Large)
Graph Tokenizer Fea. Hallu. Acc. Fea. Hallu. Acc. Fea. Hallu. Acc. Fea. Hallu. Acc.
Edge List 0.914 0.086 0.902 0.788 0.192 0.720 0.994 0.004 0.522 0.870 0.128 0.000
Adjacency List 0.898 0.102 0.894 0.808 0.166 0.774 0.990 0.006 0.288 0.918 0.082 0.000
Adjacency Matrix 0.842 0.158 0.780 0.546 0.436 0.448 0.940 0.058 0.318 0.806 0.012 0.016
Table 8: Performance comparison for GPT-4o on real-world and synthetic graphs.
Tasks Neighbor (Small) Neighbor (Large) Distance (Small) Distance (Large)
Graph Type Fea. Hallu. Acc. Fea. Hallu. Acc. Fea. Hallu. Acc. Fea. Hallu. Acc.
Real-world Graphs 0.838 0.088 0.778 0.740 0.140 0.626 0.922 0.078 0.904 0.760 0.224 0.680
Erdős-Rényi Graphs 0.996 0.004 0.986 0.984 0.016 0.730 0.980 0.020 0.946 0.864 0.146 0.702
@@
Table 18: An example of the Traveling Salesman Problem.